# Marginalizing and Conditioning Gaussians onto Linear Approximations of Smooth Manifolds with Applications in Robotics

### v1

@@ IV-A General Manifolds
Before confining to linear manifolds, we first clarify our definitions for marginalizing and conditioning onto smooth, potentially nonlinear manifolds, in preparation for Section V. Suppose we have a random variable, $\mathbf{x}\in\mathbb{R}^{n}$, an associated probability distribution, $p(\mathbf{x})$, and a manifold, $\mathcal{M}:\mathbf{f}(\mathbf{x})=\mathbf{0}$, where $\mathbf{f}(\cdot)$ is differentiable. Suppose the manifold is $m$-dimensional. That is, $\mathbf{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$.
For marginalization, suppose $\mathcal{M}$ has a projection method (i.e., a norm is defined), $\mathbf{g}_{\mathcal{M}}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, where $\mathbf{g}_{\mathcal{M}}(\mathbf{x})=\tilde{\mathbf{x}}\in\mathcal{M}\;\forall\mathbf{x}\in\mathbb{R}^{n}$. A common projection method is the closest on-manifold point measured via the Euclidean distance: $\mathbf{g}_{\mathcal{M}}(\mathbf{x})=\operatorname*{\arg\!\min}_{\mathbf{p}\in\mathcal{M}}||\mathbf{p}-\mathbf{x}||^{2}$. Then, $p_{\text{marg}:\mathcal{M}}(\mathbf{x})$ is the result of projecting $p(\mathbf{x})$ onto $\mathcal{M}$ with $\mathbf{g}_{\mathcal{M}}$. For conditioning, we simply condition $p(\mathbf{x})$ over the points on the manifold: ${p}_{\text{cond}:\mathcal{M}}(\mathbf{x})=p(\mathbf{x}|\mathbf{f}(\mathbf{x})=\mathbf{0})$. Instead of slicing and renormalizing $p(\mathbf{x})$ over fixed values, we renormalize over the slice defined by the manifold. Note that for both marginalization and conditioning, the density is zero at all points off the manifold.
@@ VI-C Covariance Extraction in Constrained GTSAM
On-manifold optimizers are prevalent in robotics but face challenges in covariance computation. This is the case for constrained GTSAM algorithms including Incremental Constrained Smoothing (ICS) [36] and Incremental Constrained Optimization (InCOpt) [4]. The original unconstrained GTSAM [5] is used for incremental smoothing and mapping. Upon receiving new information, GTSAM modifies the affected factors in the factor graph, essentially updating the batch information matrix. Thus, the unconstrained covariance can be computed as a byproduct of solving for the mean [6]. Meanwhile, constrained GTSAM is used in the presence of known hard constraints, but no longer offers the capability to compute covariances. This section restores this capability by conditioning the unconstrained covariance onto the linearized constraints. Similar to Section VI-B, we again apply (II), where $\bm{\Sigma}^{-1}$ is the unconstrained batch information matrix and $\mathbf{N}$ spans the nullspace the linearized constraints. Both $\bm{\Sigma}^{-1}$ and the linearized constraints ($\mathbf{S}$ in Table II) are already being maintained while computing the state mean. Thus, $\bm{\Sigma}_{\text{cond}}$ can be solved as a byproduct. If desired, a sparse solver can be used to efficiently compute (II).
@@
Figure 5: Visualization of the InCOpt (constrained GTSAM) output. The arrows represent the box’s end poses. Throughout the trajectory, the box is constrained to be in contact with the circular probe. The $3\sigma$ bounds of the estimated covariance form a narrow ellipse.
@@ VII Conclusion
There are many avenues for future work. More accurate techniques could be developed for marginalizing and conditioning onto nonlinear manifolds, possibly with sigma points [38][39]. For applications, covariance extraction in constrained settings has received limited attention to date, and perhaps this paper could stimulate further exploration.

### v3

@@ IV-A General Manifolds
Before confining to linear manifolds, we first clarify our definitions for marginalizing and conditioning onto smooth, potentially nonlinear manifolds, in preparation for Section V. Suppose we have a random variable, $\mathbf{x}\in\mathbb{R}^{n}$, an associated probability distribution, $p(\mathbf{x})$, and a manifold, $\mathcal{M}:\mathbf{f}(\mathbf{x})=\mathbf{0}$, where $\mathbf{f}(\cdot)$ is differentiable. Suppose the manifold is $m$-dimensional. That is, $\mathbf{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n-m}$.
For marginalization, suppose $\mathcal{M}$ has a projection method (i.e., a norm is defined), $\mathbf{g}_{\mathcal{M}}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, where for any $\mathbf{x}\in\mathbb{R}^{n}$, $\mathbf{g}_{\mathcal{M}}(\mathbf{x})=\tilde{\mathbf{x}}\in\mathcal{M}$, meaning that $\mathbf{f}(\tilde{\mathbf{x}})=\mathbf{0}$. A common projection method is the closest on-manifold point measured via the Euclidean distance: $\mathbf{g}_{\mathcal{M}}(\mathbf{x})=\operatorname*{\arg\!\min}_{\mathbf{p}\in\mathcal{M}}||\mathbf{p}-\mathbf{x}||^{2}$. Then, $p_{\text{marg}:\mathcal{M}}(\mathbf{x})$ is the result of projecting $p(\mathbf{x})$ onto $\mathcal{M}$ with $\mathbf{g}_{\mathcal{M}}$. For conditioning, we simply condition $p(\mathbf{x})$ over the points on the manifold: ${p}_{\text{cond}:\mathcal{M}}(\mathbf{x})=p(\mathbf{x}|\mathbf{f}(\mathbf{x})=\mathbf{0})$. Instead of slicing and renormalizing $p(\mathbf{x})$ over fixed values, we renormalize over the slice defined by the manifold. Note that for both marginalization and conditioning, the density is zero at all points off the manifold.
@@ VI-C Covariance Extraction in Constrained GTSAM
On-manifold optimizers are prevalent in robotics but face challenges in covariance computation. This is the case for constrained GTSAM algorithms including Incremental Constrained Smoothing (ICS) [36] and Incremental Constrained Optimization (InCOpt) [4]. The original unconstrained GTSAM [5] is used for incremental smoothing and mapping. Upon receiving new information, GTSAM modifies the affected factors in the factor graph, essentially updating the batch information matrix. Thus, the unconstrained covariance can be computed as a byproduct of solving for the mean [6]. Meanwhile, constrained GTSAM is used in the presence of known hard constraints, but no longer offers the capability to compute covariances. This section restores this capability by conditioning the unconstrained covariance onto the linearized constraints. Similar to Section VI-B, we again apply (II), where $\bm{\Sigma}^{-1}$ is the unconstrained batch information matrix and $\mathbf{N}$ spans the nullspace of $\mathbf{S}$, the linearized constraints. Both $\bm{\Sigma}^{-1}$ and $\mathbf{S}$ are already being maintained while computing the state mean. Thus, $\bm{\Sigma}_{\text{cond}}$ can be solved as a byproduct. If desired, a sparse solver similar to the one used in RCKL-SLAM can be used to efficiently compute (II).
@@
Figure 5: Visualization of the InCOpt (constrained GTSAM) output. The arrows represent the box’s end poses. Throughout the trajectory, the box is constrained to be in contact with the circular probe. The $3\sigma$ bounds of the estimated covariance form a narrow (i.e., eccentricity is close to 1) ellipse.
@@ VII Conclusion
For future work, more accurate techniques could be developed for marginalizing and conditioning onto nonlinear manifolds, possibly with sigma points [38][39]. Covariance extraction in applications with constrained settings has received limited attention to date, and perhaps this paper could stimulate further exploration.